OpenStudy (jb1625):

Which expressions, when evaluated with y = –7, are positive in value? Choose all answers that are correct. A. 18 + 2y B. –3y + (–17) C. –2y – 23 D. –30 – 4y

OpenStudy (yoloshroom):

18+2(-7) = 18-14 = 4. - Positive -3(-7)-17=21-17=4 - Positive -2(-7)-23=-14-23=-37 - Negative -30 -4(-7)=-30+28 = -2 - Negative So Both A and B would be Positive, and C and D would be negative right?

OpenStudy (jb1625):

So the answer would be A. and B.?

OpenStudy (yoloshroom):

I believe so :D

OpenStudy (jb1625):

Thank you :D

OpenStudy (yoloshroom):

No problem m8 :)

OpenStudy (jb1625):

What does that mean I have always wondered ?

OpenStudy (yoloshroom):

m8 = Mate ^_^ just like saying "Buddy" or something along those lines

OpenStudy (jb1625):

Oh thx